OpenStudy (anonymous):

Find the LCM of y^2+8y+16 and y^2+y-12

OpenStudy (anonymous):

factoring these polynomials we obtain:\[y^2+8y+16=(y+4)^2\]\[y^2+y-12=(y+4)(y-3)\]So their GCD is (y + 4). Now using the formula:\[LCM(a,b)=\frac{a\cdot b}{GCD(a,b)}\]i'll multiply the two polynomials and divide by their GCD. To make things easier, i'll use the factored form:\[LCM=\frac{(y+4)^2\cdot (y+4)(y-3)}{(y+4)}=(y+4)^2 (y-3)=y^3+5y^2-8y-48\]

hero (hero):

Gee Whiz. Such an in depth explanation

OpenStudy (anonymous):

I feel that getting the GCD of two objects is easier than getting the LCM. Thats why i do it this way. Its "plug n' chug" after you find the GCD.